# Identifying Uncertainty in Self-Adaptive Robotics with Large Language Models

### v1

@@ 1 Introduction
A key challenge in self-adaptive robotics is managing uncertainty throughout the entire engineering lifecycle, from initial design to active operations [3]. Uncertainty may arise from various sources, such as unpredictable environmental conditions, sensor and actuator noise, and human-robot and robot-robot interactions. Such factors directly influence a robot’s dependability, ultimately impacting its overall performance and decision-making capabilities. A crucial step in managing uncertainty within these systems is identifying the sources and potential impacts of uncertainties at early stages of the robotics software engineering lifecycle [3]. However, it is challenging due to the inherent complexity of self-adaptive robots, including their complex interactions, evolving configurations, and adaptive behaviors, as well as the lack of comprehensive knowledge about unpredictable operational contexts and environmental dynamics. Consequently, practitioners often depend on their intuition and insights from previous experiences with similar systems to identify and manage uncertainties in self-adaptive robotics.
@@
The growing adoption of large language models (LLMs) with advanced text capabilities, such as generation, analysis, and comprehension [4], extends to various fields, including software engineering [5, 6]. In this article, we explore LLMs’ ability to assist practitioners in identifying uncertainties when developing self-adaptive robots throughout the entire software engineering lifecycle, aiming to provide a more systematic and comprehensive approach that improves over the current practice of relying primarily on intuition and experience. To achieve this, we provided LLMs with robots’ software requirements and carefully designed prompts containing clear instructions and an uncertainty questionnaire. After obtaining and compiling responses from the LLMs, we conducted in-person sessions with the practitioners to discuss and gather their feedback. Using input provided by practitioners, we analyzed the quality, relevance, and usefulness of the uncertainties identified by LLMs.
@@
For evaluation, we selected 10 advanced LLMs with varying capabilities: Gemini Pro 1.5, Perplexity Sonar, Nemotron 70B, Nova Pro, Mistral Large 2411, LLama 3.3 70B, o1 Preview, GPT-4o, Gemini Flash 2.0, and Claude 3.5 Sonnet. We evaluated these LLMs across four real-world industrial robotics use cases: Industrial Disassembly Robot (IDR), Warehouse Robotic Swarm (WRS), Prolonged Hull of an Autonomous Vessel (PHAV), and Human-Robotic Interaction (HRI). To evaluate the LLM-generated responses regarding uncertainties, we collected feedback from the practitioners involved in each case. Results demonstrated that practitioners agreed with 65–72% of LLMs-generated responses in the IDR case, 83–88% in the WRS case, 79–86% in the PHAV case, and 63–76% in the HRI case. In addition, insights from discussions with practitioners suggested that LLMs have the potential to comprehend robots’ software requirements, provide logical responses to uncertainties, identify novel uncertainties, and demonstrate their usefulness in practical contexts.
@@ 3 Evaluation Methodology
Goal: Our goal is to evaluate the effectiveness of LLMs in identifying uncertainties in self-adaptive robotics and to determine their practical value for the practitioners.
@@
Step : We collected documented robotics software requirements from the practitioners. These documents included detailed descriptions of each use case scenario, clearly specifying the operational context, expected behavior, objectives, and constraints. Moreover, they explicitly outlined hardware characteristics, including details of sensors and actuators, as well as thorough software requirements defining the system architecture, functionalities, and performance expectations.
@@
Steps : The subsequent steps, from four to six, involve interaction with LLMs. In step four, we perform prompt engineering to prepare prompts for LLMs. Specifically, we used the role-prompting technique, which asks the model to assume a particular role or expertise. We provided LLMs with the requirement documents and asked them to assume the role of expert analyst. Next, we provided step-by-step instructions to follow, which included carefully reading and comprehending the requirements document, utilizing their robotics knowledge, understanding the purpose of the questionnaire, and responding to questions concisely with brief justifications. After finalizing the prompt (available on GitHub [^1] https://github.com/Simula-COMPLEX/Prompt4RoboUI), the subsequent steps involve iteratively prompting all selected LLMs for each case study and collecting and compiling LLMs’ responses. Once we have obtained responses from all LLMs and case studies, the next step is to prepare confirmation questionnaires for each partner.
@@
Steps : In the eighth step, we conducted four individual in-person sessions during project meetings with the practitioners involved in each case study. Each session included one senior and one junior participant from the respective case study. Each session lasted 40 minutes and followed a consistent format. At the beginning of each session, we provided participants with a concise overview of the planned activities and time allocation, presented the revised uncertainty questionnaire, and explained the objectives and motivations of the activity. After this introductory session of 10 minutes, we distributed the confirmation questionnaire to the participants. We allocated 20 minutes for the questionnaire and 10 minutes of verbal discussion. During the questionnaire activity, participants were asked to assess each response and indicate their level of agreement based on their experience and knowledge of the case study. In the ninth step, we collected participants’ feedback on the confirmation questionnaire through Google Forms. During the verbal discussion, we gathered participants’ opinions on the LLM-generated responses, the LLMs’ understanding of the robotic case study, and their potential for identifying uncertainties. For this discussion, we opted to take informal notes instead of formally recording the feedback.
@@ 4.1 IDR Results
Figure 2(a) shows that the responses generated by GPT-4o received the highest number of agreements and the lowest number of disagreements. The second highest number of agreements was observed for the responses generated by o1 Preview and LLama 3.3 70B. Moreover, practitioners indicated neutral agreement only for the responses generated by Nemotron 70B and Nova Pro, at 6% and 5%, respectively. Regarding disagreements, the responses generated by Gemini Pro 1.5, Mistral Large 2411, and Claude 3.5 Sonnet received a notably higher number of disagreements compared to other models. Among these, practitioners strongly disagreed with most of the responses obtained from Gemini Pro 1.5 and Mistral Large 2411. For simple disagreement opinions, practitioners showed a high level of disagreement with responses generated by LLama 3.3 70B and Claude 3.5 Sonnet at 18%, whereas responses from GPT-4o received the lowest level of disagreement at 11%. Overall, the results indicate that practitioners agreed with more than 60% of the LLM-generated responses, including a majority of strong agreements.
@@ 4.2 WRS Results
From Figure 2(b), it can be observed that practitioners showed a high number of strong agreements with responses generated by Claude 3.5 Sonnet. For general agreements (including agreeing and strongly agreeing), practitioners agreed with the responses generated by Nemotron 70B, Nova Pro, and o1 Preview. Interestingly, none of the practitioners strongly disagreed with any of the responses. Moreover, overall disagreements are noticeable but remain up to 11%. For the responses from Gemini Flash 2.0, the number of disagreements is low compared to other models. The percentage of neutral agreements is 8% for the responses obtained from Gemini Pro 1.5, Perplexity Sonar, and LLama 3.3 70B. The overall results show that the practitioners agreed with more than 80% of the LLM-generated responses, indicating the high practicality of LLMs in this case study.
@@ 4.3 PHAV Results
Results in Figure 2(c) indicate that the responses generated by LLama 3.3 70B received the highest level of agreement and the fewest disagreements. Following closely, the responses generated by Gemini Flash 2.0 received the second-highest score in terms of agreement. For LLama 3.3 70B and Gemini Flash 2.0, practitioners did not indicate any neutral agreements. In contrast, other models exhibited neutral agreements, with practitioners showing up to 4% and 3% for responses generated by Mistral Large 2411 and o1 Preview, respectively. Furthermore, the number of strong disagreements remains relatively low, with 3% for the responses from Mistral Large 2411, LLama 3.3 70B, and o1 Preview, and 4% for other models. Similarly, the highest number of disagreements were observed for responses generated by Gemini Pro 1.5, while Mistral Large 2411, LLama 3.3 70B, and o1 Preview had the lowest disagreement rates. In general, practitioners expressed positive opinions for more than 80% of the LLM-generated responses, highlighting the high effectiveness of LLMs.
@@ 4.4 HRI Results
Figure 2(d) shows that the practitioners strongly agreed with many of the responses (48%) generated by Perplexity Sonar and Mistral Large 2411, followed by GPT-4o with 47% strong agreements. Notably, practitioners demonstrated strong agreement with over 40% of all LLMs’ responses. Regarding the overall agreement, the highest number of positive opinions were particularly observed for responses generated by Nova Pro and o1 Preview. For neutral agreements, practitioners indicated the highest neutrality for responses generated by Perplexity Sonar, while Nova Pro received the least. In disagreements case, practitioners strongly disagreed with the smallest percentage (4–5%) of LLM-generated responses. In contrast, the percentage of simple disagreements ranged from 15% to 23%. Among all models, the responses generated by Perplexity Sonar and LLama 3.3 70B had the highest number of disagreements, while fewest disagreements were observed for Gemini Pro 1.5. Overall, practitioners provided positive opinions for more than 70% of the responses generated by LLMs, demonstrating their usefulness for this case study.
@@ 4.5 Insights
Below, we present key insights from results analysis and discussions with practitioners.
@@ 4.5.1 LLMs’ Familiarity with Domain
When we asked practitioners from all case studies whether they believed LLMs could effectively understand the case study specifications and domain, all participants agreed that LLMs’ responses indicated a strong grasp of the case study. This highlights LLMs’ ability to comprehend text by using the provided information about various robots and their prior knowledge from related domains.
@@ 4.5.2 Logical Responses
4.5.3 Novel Uncertainties
We asked practitioners whether they found the uncertainties identified by LLMs to be realistic and whether any of them were previously unknown yet logically valid. They unanimously agreed that the LLM-identified uncertainties were realistic and were particularly surprised by some previously unconsidered uncertainties, which they found valuable.
@@ 4.5.4 Knowledge Gaps in Engineering Phases
During the discussion, practitioners revealed a lack of knowledge about uncertainties across different robotics software engineering phases. Some participants were involved exclusively in development, others in testing, and some solely in the operations phase. Those involved in the development phase had limited awareness of uncertainties that could arise during operations, while those in the operations phase were unfamiliar with uncertainties in the design phase. Consequently, they could not confirm uncertainties outside their expertise and either disagreed or remained neutral when evaluating LLM-identified uncertainties related to phases they were not directly involved in.
4.5.5 Uncertainty in Engineering Phases
Although all practitioners agreed that uncertainties arise in every engineering phase, the majority emphasized that the most uncertainties are typically encountered during the testing phase. Those involved in the operational phase highlighted that uncertainties are equally prevalent during live operations. This was especially noticed in the autonomous vessel case study, where real-world operations in complex environments often introduce significant environmental uncertainties that may go undetected during simulation-based testing. Therefore, depending on the complexity and dynamics of the environment, both the testing and operational phases are likely to encounter uncertainties. Furthermore, all practitioners unanimously confirmed that environmental dynamics are a major source of uncertainty.
@@ 4.5.6 Managing Uncertainty
4.5.7 LLM Recommendations
While all LLMs exhibited varying levels of usefulness in identifying uncertainties, their responses differed in depth and detail. Models such as Claude 3.5 Sonnet generated concise responses that lacked comprehensive explanations, whereas others, like o1 Preview, provided more detailed and well-reasoned answers for each point. This was particularly notable when LLMs were asked to provide example uncertainty scenarios for specific robotic cases, which practitioners endorsed and found interesting. Thus, based on the results analysis, we recommend LLama 3.3 70B, o1 Preview, GPT-4o, Mistral Large 2411, and Nova Pro as the most effective LLMs for uncertainty analysis.
@@ 4.6 Threats to Validity
A potential threat to external validity is the limited sample size of eight practitioners in four use cases. Furthermore, the robotic cases studied specifically focused on small subsystems dedicated to self-adaptive robotic behaviors. Therefore, results might not generalize to all robotic systems, which is a common threat in empirical studies. An internal validity threat may occur due to participants’ personal biases and familiarity with robotic systems. Although our sessions with practitioners involved two participants per case (one junior and one senior), their backgrounds and confidence levels may have influenced their responses. To handle this, we included open discussions during sessions to clarify interpretations. A potential conclusion validity threat may occur due to variability in practitioners’ experience levels and the subjective nature of agreement measurements (i.e., Likert levels). To manage this, we used standardized procedures in all sessions, providing clear instructions and consistent timing, although subjective assessment remains an inherent challenge. Moreover, to address the validity threat from informal notes during verbal discussions, we sent the draft article to all participating practitioners for review and revised it based on their feedback. A possible threat involves the clarity of the uncertainty questionnaire and confirmation questionnaire. To mitigate this, we developed the uncertainty questionnaire based on existing frameworks and revised it based on practitioners’ feedback. We also used the standard Likert scale to capture agreement levels for the confirmation questionnaire precisely.

### v2

@@ 1 Introduction
A key challenge in self-adaptive robotics is managing uncertainty throughout the entire engineering lifecycle, from initial design to active operations [3]. Uncertainty may arise from various sources, such as unpredictable environmental conditions, sensor and actuator noise, and human-robot and robot-robot interactions. Such factors directly influence a robot’s dependability, ultimately impacting its overall performance and decision-making capabilities. A crucial step in managing uncertainty within these systems is identifying the sources and potential impacts of uncertainties at early stages of the robotic software engineering lifecycle [3]. However, it is challenging due to the inherent complexity of self-adaptive robots, including their complex interactions, evolving configurations, and adaptive behaviors, as well as the lack of comprehensive knowledge about unpredictable operational contexts and environmental dynamics. Consequently, practitioners often depend on their intuition and insights from previous experiences with similar systems to identify and manage uncertainties in self-adaptive robotics.
@@
The growing adoption of large language models (LLMs) with advanced text capabilities, such as generation, analysis, and comprehension [4], extends to various fields, including software engineering [5]. In this article, we explore LLMs’ ability to assist practitioners in identifying uncertainties when developing self-adaptive robots throughout the entire software engineering lifecycle, aiming to provide a more systematic and comprehensive approach that improves over the current practice of relying primarily on intuition and experience.
@@
For evaluation, we selected 10 advanced LLMs with varying capabilities: Gemini Pro 1.5, Perplexity Sonar, Nemotron 70B, Nova Pro, Mistral Large 2411, LLama 3.3 70B, o1 Preview, GPT-4o, Gemini Flash 2.0, and Claude 3.5 Sonnet. We evaluated these LLMs across four real-world industrial robotic use cases: Industrial Disassembly Robot (IDR), Warehouse Robotic Swarm (WRS), Prolonged Hull of an Autonomous Vessel (PHAV), and Human-Robotic Interaction (HRI). To evaluate the LLM-generated responses regarding uncertainties, we conducted a study with practitioners from each case to collect their feedback. Results demonstrated that practitioners agreed with 65–72% of LLMs-generated responses in the IDR case, 83–88% in the WRS case, 79–86% in the PHAV case, and 63–76% in the HRI case. In addition, insights from discussions with practitioners suggested that LLMs have the potential to comprehend robots’ software requirements, provide logical responses to uncertainties, identify overlooked uncertainties, and demonstrate their usefulness in practical contexts.
This work evaluates LLMs’ potential to support practitioners in identifying uncertainties in self-adaptive robotics. We focus on system-level aspects rather than exploring inter-module interactions or specific uncertainty categories. Within this scope, the main contributions are: (i) a novel LLM-prompting method for uncertainty identification based on robotic requirements, (ii) an LLM evaluation study with practitioners from four different robotic industries, and (iii) actionable insights and future directions for researchers and practitioners.
@@ 3 Evaluation Methodology
Goal: Evaluate LLMs’ effectiveness in identifying uncertainties in self-adaptive robotics and to determine their value for practitioners.
@@
Step : We collected documented robotic software requirements from the practitioners. These documents included detailed descriptions of each use case scenario, clearly specifying the operational context, expected behavior, objectives, and constraints. Moreover, they explicitly outlined hardware characteristics, including details of sensors and actuators, as well as thorough software requirements defining the system architecture, functionalities, and performance expectations.
@@
Steps : The subsequent steps, from four to six, involve interaction with LLMs. In step four, we performed prompt engineering to prepare prompts for LLMs. We initially tried zero-shot prompting with direct instructions, but the responses lacked precision. As our focus is on supporting experts in identifying uncertainties using LLMs, we adopted role-based prompting after exploring zero-shot and self-consistency prompting. Role-based prompting lets LLMs assume a specific role (e.g., engineer or tester) to give task-specific responses. Therefore, we developed a role-based prompt (available on Zenodo [7]) with step-by-step instructions and the uncertainty questionnaire. The instructions include reading requirements, applying robotic knowledge, understanding the questionnaire’s purpose, and providing concise, justified responses. The next steps involve iteratively prompting all LLMs for each case study and compiling their responses. The prompting goal is to obtain answers to uncertainty questions based on their understanding of the robotic context. For this, we provided each LLM with the requirement document and the prompt, then collected their responses. To mitigate hallucination risks, we used self-consistency prompting by re-prompting the LLMs with the same prompt three times (and up to five times in cases of variability) and selecting the most consistent response. After each prompting iteration, final responses were stored and labeled with the corresponding LLM.
@@
Steps : In the eighth step, we conducted four individual in-person sessions during project meetings with the practitioners involved in each case study. Each session included one senior and one junior participant from the respective case study. Each session lasted 40 minutes and followed a consistent format. At the beginning of each session, we provided participants with a concise overview of the planned activities and time allocation, presented the revised uncertainty questionnaire, and explained the objectives and motivations of the activity. Following the 10-minute introductory session, we distributed the confirmation questionnaire to the participants via Google Forms. We allocated 20 minutes for the questionnaire and 10 minutes for verbal discussion. During the questionnaire activity, participants were asked to assess each response to the uncertainty questionnaire and indicate their level of agreement based on their experience and knowledge of the case study. The confirmation questionnaire included sections for each uncertainty question, showing LLM responses and a Likert scale for agreement. Participants had sufficient time to review and assess responses thoughtfully. They submitted their assessments using Google Forms. This was followed by a verbal discussion session to gather broader feedback without revisiting the submitted responses. During the verbal discussion, we gathered participants’ opinions on the LLM-generated responses, the LLMs’ understanding of the robotic case study, and their potential for identifying uncertainties. For this discussion, we opted to take informal notes instead of formally recording the feedback.
@@ 4.1 IDR Results
Figure 2(a) shows that GPT-4o’s responses received the highest number of agreements and the lowest disagreements, followed by o1 Preview and LLama 3.3 70B. Neutral agreements were reported for only Nemotron 70B (6%) and Nova Pro (5%). Gemini Pro 1.5, Mistral Large 2411, and Claude 3.5 Sonnet had the highest levels of disagreement. Overall, practitioners agreed with over 60% of the LLM responses, mostly with strong agreements.
@@ 4.2 WRS Results
Figure 2(b) demonstrates that practitioners strongly agreed most often with responses generated by Claude 3.5 Sonnet, while overall agreements were high for Nemotron 70B, Nova Pro, and o1 Preview. No strong disagreements were reported, with overall disagreements low at 11%. Moreover, neutral agreements are 8% for responses from Gemini Pro 1.5, Perplexity Sonar, and LLama 3.3 70B. In general, practitioners agreed with over 80% of the LLM responses, highlighting LLMs’ high practicality in this case.
@@ 4.3 PHAV Results
Figure 2(c) indicates that the responses generated by LLama 3.3 70B had the highest agreement and lowest disagreements, followed by Gemini Flash 2.0. No neutral agreements were reported for these two models, while Mistral Large 2411 and o1 Preview had 4% and 3%, respectively. Furthermore, strong disagreements were low overall, while highest for Gemini Pro 1.5. Overall, practitioners expressed positive opinions for over 80% of the LLM responses, highlighting LLMs’ high effectiveness.
@@ 4.4 HRI Results
Figure 2(d) shows that practitioners strongly agreed with over 40% of responses from all LLMs, with Perplexity Sonar, Mistral Large 2411, and GPT-4o reaching 48%, while Nova Pro and o1 Preview received the highest overall agreement. For neutral agreements, Perplexity Sonar had the most and Nova Pro the least. Moreover, practitioners strongly disagreed with only 4–5% of LLM responses, while simple disagreements ranged from 15–23%, with Nemotron 70B and LLama 3.3 70B having the most disagreements and Gemini Pro 1.5 the lowest. In general, practitioners agreed with over 70% of LLM responses, demonstrating LLMs’ usefulness in this case.
4.5 Reliability and Confidence Analysis
First, we analyzed the inter-rater reliability using Cohen’s Kappa, which revealed that both participants unanimously agreed or disagreed in each case, resulting in no false positives or false negatives (detailed results are provided in Tables 1, 2, 3 and 4 and discussed in Appendix A). Next, we analyzed 95% confidence intervals using the bootstrap method, which demonstrated high confidence in the results in all case studies (detailed results are presented in Tables 5, 6, 7 and 8 and further discussed in Appendix B). Overall, the strong agreement levels, combined with minimal variation in disagreement and neutrality, demonstrate that the results for each case study are reliable and consistent. Detailed findings are also available in the online repository [7].
@@ 4.6 Insights
We present key insights from results analysis and discussions with practitioners.
@@ 4.6.1 LLMs’ Familiarity with Domain
When we asked practitioners from all case studies whether they believed LLMs could effectively understand the case study specifications and domain, all participants agreed that LLMs’ responses indicated a strong grasp of the case study. This highlights LLMs’ ability to comprehend text using the provided information about various robots and their prior knowledge from related domains.
@@ 4.6.2 Logical Responses
4.6.3 Overlooked Uncertainties
We asked practitioners whether they found the uncertainties identified by LLMs to be realistic and whether any of them were previously unknown yet logically valid. They unanimously agreed that the LLM-identified uncertainties were realistic and were particularly surprised by some previously unconsidered uncertainties, which they found valuable. Below, we present two uncertainty examples for each case study, classified by type, source, lifecycle stage, and alignment with existing taxonomies. In IDR, variations in laptop components can be screws (aleatoric; hardware; operational; corresponds [6]) and cable routing issues during disassembly (epistemic; internal; operational; corresponds [8]). In WRS, LiDAR sensor occlusion may disrupt robot navigation (aleatoric; environmental; operational; corresponds [6]) and mislocalize due to odometry errors (epistemic; software/internal; operational; corresponds [8]). In PHAV, uncertainty can result from transferring models between vessels with differing dynamics (epistemic; software/internal; testing; corresponds [8]) and unexpected weather changes with complex wave-current interactions (aleatoric; environmental; operational; corresponds [6]). In HRI, unnecessary interactions with humans and objects may cause frequent re-planning (aleatoric; external; operational; corresponds [9]) and variations in the workspace setup or additional obstacles (epistemic; environmental/external; operational; corresponds [8]).
@@ 4.6.4 Knowledge Gaps in Engineering Phases
Practitioners revealed a lack of knowledge about uncertainties across different robotic software engineering phases. Some participants focused only on development, others on testing, or operations. Developers had limited awareness of operational uncertainties, while those in the operations phase were unfamiliar with design-phase uncertainties. Consequently, they could not confirm uncertainties outside their expertise and either disagreed or remained neutral when evaluating LLM-identified uncertainties related to phases they were not directly involved.
4.6.5 Uncertainty in Engineering Phases:
Although all practitioners agreed that uncertainties arise in every engineering phase, the majority emphasized that the most uncertainties are typically encountered during the testing phase. Those involved in the operational phase highlighted that uncertainties are equally prevalent during live operations. This was especially noticeable in the autonomous vessel case study, where real-world operations in complex environments often introduce significant environmental uncertainties that may go undetected during simulation-based testing. Therefore, depending on the complexity and dynamics of the environment, both the testing and operational phases are likely to encounter uncertainties. Furthermore, all practitioners unanimously confirmed that environmental dynamics are a major source of uncertainty.
@@ 4.6.6 Managing Uncertainty
4.6.7 LLM Uncertainty Identification Trends
LLMs effectively identify uncertainties from detailed hardware and software specifications, as evident from the WRS results. However, LLMs often apply their general uncertainty knowledge to all aspects mentioned in the requirements, which may not fit specific systems. For example, uncertainties from human error are less relevant for autonomous vessels (PHAVs) due to experienced operators and well-defined protocols, though such uncertainties may arise in WRS. Regarding uncertainty occurrence, uncertainties arise primarily from hardware issues and dynamic environments, followed by ML models in MAPLE-K.
4.6.8 LLM Reasoning
While we observed no major differences among LLMs with varying reasoning capabilities, reasoning models like o1 Preview generated more detailed and well-structured responses, whereas non-reasoning models like Perplexity Sonar still attempted to incorporate some level of reasoning in their outputs. Furthermore, general-purpose models like GPT-4o often generated broader, less context-specific reasoning, which frequently led to practitioner disagreement. In particular, the example uncertainty scenarios generated by the reasoning models were interesting and endorsed by practitioners. Therefore, we recommend using reasoning models with advanced prompting techniques for effective uncertainty analysis.
4.6.9 Actionable Recommendations
Given LLMs’ varied reasoning capabilities, practitioners shall adopt a combination of prompting strategies, such as few-shot prompting for providing examples and chain-of-thought prompting for step-by-step reasoning, to iteratively refine responses and enhance the identification of true uncertainties. To minimize hallucination risks, practitioners identifying uncertainties at the design phase should communicate LLM responses with teams across engineering phases, such as operations and testing, and incorporate their feedback into subsequent prompts. As best practices, practitioners should use a second LLM to cross-check responses for alternative perspectives and greater confidence in the results, and remove sensitive information from documents provided to LLMs.
@@ 4.7 Threats to Validity
A potential threat to external validity is the limited sample size of eight practitioners involved in four robotic cases focused on self-adaptive behaviors. However, involving more participants was infeasible due to a limited number of personnel allocated to the project, so we included those with sufficient experience and familiarity with robotics to address this. Although the results may not generalize to all robotic systems, such limitations are common in empirical studies involving industry collaborations. An internal validity threat may occur due to participants’ personal biases and familiarity with robotic systems. Although our sessions with practitioners involved two participants per case (one junior and one senior), their backgrounds and confidence levels may have influenced their responses. To handle this, we included open discussions during sessions to clarify interpretations. A potential conclusion validity threat may occur due to variability in practitioners’ experience levels and the subjective nature of agreement measurements (i.e., Likert levels). To mitigate this, we used standardized procedures in all sessions, providing clear instructions and consistent timing, although subjective assessment remains an inherent challenge. Moreover, we used Cohen’s Kappa to analyze inter-rater reliability and bootstrap confidence intervals to evaluate the confidence in the results. A possible threat involves the clarity of the uncertainty questionnaire and confirmation questionnaire. To manage this, we developed the uncertainty questionnaire based on existing frameworks and revised it based on practitioners’ feedback. We also used the standard Likert scale to capture agreement levels for the confirmation questionnaire precisely.